OpenStudy (tray):

PLEASE HELP ILL MEDAL A scuba diver was exploring a reef 32.12 m below sea level. The diver ascended to the surface at a rate of 8.8 m/min. (a) Write an equation to model the situation. Use the formula d = rt. (b) Explain why it is reasonable to write the distance, d, as a positive number. (c) How long did it take the diver to reach the surface? Answer a b then c

OpenStudy (jalil.h):

d = 32.12 m; r = 8.8 m / min; t = ? The formula is: d = r t t = d / r t = 32.12 m : 8.8 m/min t = 3.65 min = 3 minutes and 39 seconds

OpenStudy (tray):

So (a) 32.12m=8.8 (b) What (c) 3.65= 3 mins and 39 seconds

OpenStudy (welshfella):

(a) d = rt plug in the given values 32.12 = 8.8t

OpenStudy (welshfella):

(b) He is ascending so its reasonable to make the distance d positive.

OpenStudy (welshfella):

Also the time and the rate will be positive so the distance will be positive

OpenStudy (tray):

(c) ???

OpenStudy (tray):

How long did it take??? (c)????

OpenStudy (mathmale):

Note: if d = r t, then t = d / r. Just divide the distance by the rate (speed).

OpenStudy (tray):

All I need is (c) PLEASEEEEEE TELLLL MEEEEEEEEEE :-(

OpenStudy (mathmale):

Please be sure to include units of measurement in your work and your work. It's important to me (if not to you?) that you understand the principles involved here. d = r t is a general formula: distance equals rate times time. You MUST know this. If you want to calculate the amount of time required for the diver to ascend at 8.8 m/min when the distance involved is 32.12 m, you must divide this distance by the rate. Give it a try, please. Share your work.

OpenStudy (tray):

3.65, I dont know

OpenStudy (mathmale):

Where are your units of measurement? Your 3.65 is OK, but 3.65 can't stand alone.

OpenStudy (mathmale):

What is\[\frac{ 32.12~m }{ 8.8\frac{ m }{ \sec } }?\]

OpenStudy (mathmale):

Notice how you may cancel "m" here. m/m = 1

OpenStudy (tray):

3.65m/sec ???

OpenStudy (mathmale):

What is \[\frac{ 1 }{ \frac{ 1 }{ \sec } }?\]

OpenStudy (tray):

1?

OpenStudy (mathmale):

But do we really measure time in meters per second?

OpenStudy (tray):

Idk

OpenStudy (mathmale):

What do you know and what do you not know? We have the opp to clear things up here. I'll guide you but will not give you a direct answer.

OpenStudy (mathmale):

Once again: You must divide the distance by the rate (speed).

OpenStudy (mathmale):

time = distance / speed

OpenStudy (mathmale):

\[time=t=\frac{ distance }{ speed}=\frac{ 32.12~m }{ 8.8\frac{ m }{ \sec } }=?\]

OpenStudy (mathmale):

You have already gotten the numeric value correct; it is 3.65. But...3.65 what? Please simplify: \[\frac{ 1 }{ \frac{ 1 }{ \sec } }\]

OpenStudy (tray):

3.65 m/sec

OpenStudy (mathmale):

which is the logical unit of measurement for time here? meters per second, or seconds, or something else?

OpenStudy (mathmale):

Does it sound right to say "The time elapsed was 5 meters per second?"

OpenStudy (mathmale):

How many hours of sleep do you get each night?

OpenStudy (tray):

It took 3.65m/sec for the diver to reach the surface

OpenStudy (tray):

8-9hours y

OpenStudy (mathmale):

But m/sec is speed, not time. You want time, not speed.

OpenStudy (mathmale):

You gave me the correct unit of measurement there: You get 9 hours of sleep. Great. correct. you would not answer with "9 miles per hour," would you?

OpenStudy (tray):

so seconds per m

OpenStudy (mathmale):

Why does 'm' show up here?

OpenStudy (tray):

3.6 minuntes?

OpenStudy (mathmale):

Note this example: 1 divided by 1/8 is\[\frac{ 1 }{ \frac{ 1 }{ 8 } }=8\]

OpenStudy (mathmale):

Yes. It takes the diver 3.65 sec to ascend 32.10 m. Yes.

OpenStudy (mathmale):

In summary, time = distance / rate.

OpenStudy (tray):

Okay so it takes the diver 3.65sec to ascend 32.10m

OpenStudy (mathmale):

that's completely correct. Very good!

OpenStudy (tray):

Thank you!!